U.S. Steel is considering a plant expansion to produce austenitic, precipitation hardened, duplex, and martensitic stainlessFYI: $7.17 OR APR $7.20 MILLION IS WRONG...
Answer #1

ANSWER:

In order to find the yearly annual revenue we will equate the pw to zero.

pw = initial cost + cost 1 year from now(p/f,i,n) + operating costs per year(p/a,i,n) + annual revenue(p/a,i,n)

0 = - 17,000,000 - 10,000,000(p/f,16%,8) - 1,300,000(p/a,16%,8) + annual revenue(p/a,16%,8)

17,000,000 = - 10,000,000 * 0.31 - 1,300,000 * 4.34 + annual revenue * 4.34

17,000,000 = - 3,050,254.57 - 5,646,668.16 + annual revenue * 4.34

17,000,000 + 3,050,254.57 + 5,646,668.16 = annual revenue * 4.34

25,696,922.73 = annual revenue * 4.34

annual revenue = 25,696,922.73 / 4.34

annual revenue = 5,916,055

so the company must make $5,916,055 annually.